OpenStudy (anonymous):

the mean average temperature of Fairbanks, Alaska, is 27 degrees f. the monthly average temperatures vary between 36.5 degrees above and below this value. If t=0 represents February, the phase shift of the sine function is 2. a)write a model for the average monthly temperate in Fairbanks, Alaska. b) according to your model, what is the average temperate in November. 1) a) y=18.3 sin(t/6*3.14+3.14/3)+27 b)y(10)=8.7 degrees 2) a) y=27 sin(t/6*3.14+2(3.14)/3)+36.5 b) y(10)=23 degrees 3) a) y=18.3 sin(t/12*3.14-3.14/6)+27 b) y(9)=44.6 degrees

OpenStudy (anonymous):

4) a)y=36.5 sin(t/6*3.14-3.14/3)+27 b) y(9)=8.8 degrees

OpenStudy (anonymous):

@agent0smith @razor99 @ranga @phi hey can you guy help me

OpenStudy (phi):

sin(x) varies from a min value of -1 up to a max value of +1 0 is the mean value. if you want your sin to vary between 36.5 degrees above and below 27º F, you would want to multiply it by 36.5 (and add a 27) so without knowing anything else, look for an answer that has 27 + 36.5 sin( stuff)

OpenStudy (anonymous):

what do u mean by stuff? @phi

OpenStudy (anonymous):

i think it's # 2?? @phi @agent0smith

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

okay but i thought the amp was the highest point which would be 24+36.5

OpenStudy (agent0smith):

Amplitude is how far it varies above or below the average.

OpenStudy (anonymous):

so it either 1 or 3

OpenStudy (anonymous):

i think is 3 actually because November is the 9th month

OpenStudy (agent0smith):

" average temperature of Fairbanks, Alaska, is 27 degrees f." this means there needs to be a +27 "the monthly average temperatures vary between 36.5 degrees above and below this value." this means the amplitude is 36.5, not half of that, since it can be 36 above the mean or 36 below the mean. So 36.5sin(stuff) + 27 Question seems to be lacking a correct answer.

OpenStudy (agent0smith):

"average temperatures vary between 36.5 degrees above and below this value. " Actually I think this means it varies by 36.5 from the max to the min. So amplitude is half that... it's poorly worded.

OpenStudy (agent0smith):

So 18.25sin(stuff) + 27

OpenStudy (anonymous):

i put in # 3 but it was actually number 4 i just summit it :/

OpenStudy (anonymous):

thanks for the help tho @agent0smith @phi

OpenStudy (agent0smith):

lol you probably should've showed #4 as an option...

OpenStudy (agent0smith):

I didn't even notice it in the second post...

OpenStudy (agent0smith):

btw screenshots work better than this too... these aren't easy to read. 1) a) y=18.3 sin(t/6*3.14+3.14/3)+27 b)y(10)=8.7 degrees 2) a) y=27 sin(t/6*3.14+2(3.14)/3)+36.5 b) y(10)=23 degrees 3) a) y=18.3 sin(t/12*3.14-3.14/6)+27 b) y(9)=44.6 degrees

OpenStudy (anonymous):

i'll do that next next lol thanks agin for the help mane @agent0smith